# Background risk model in presence of heavy tails under dependence

### v2

@@ 3. Application on bi-dimensional renewal risk model
where by * we can take either $\max$ or and by $T_{*}$ we consider the
@@
We can observe that in this risk model, except the dependence structures appearing in the conditions on the claims, the two business lines are also dependent through the common renewal process.
@@
$\leq\sum_{n=m+1}^{\left\lfloor(x\wedge y)/L\right\rfloor}n^{2}\,\mathbf{P}\left[\sum_{i=1}^{n}X_{i}>\dfrac{x}{n}\,,\;\sum_{j=1}^{n}Y_{j}>\dfrac{y}{n}\right]\,\mathbf{P}[N(T)=n]+\mathbf{P}\left[N(T)>\dfrac{x\wedge y}{L}\right]$
@@
$\leq K\,\overline{F}\left(x\right)\,\overline{G}\left(y\right)\,\mathbf{E}\left[[N(T)]^{2\,(p+1)}{\bf 1}_{\{N(T)>m\}}\right]\,,$
@@
$\qquad\qquad\leq\dfrac{K}{M}\lim_{m\to\infty}\dfrac{1}{\lambda(T)}\,\mathbf{E}\left[[N(T)]^{2\,(p+1)}{\bf 1}_{\{N(T)>m\}}\right]=0\,,$ (3.17)
where the constant $M>0$, stems from relation (3.16), while in the last step we used [46, Lem. 3.2].
Now, we estimate the $I_{1}(x,\,y;\,T)$. At first, for the inter-arrival times $\{\theta_{1},\,\ldots,\,\theta_{n+1}\}$, through the dominated convergence theorem, by Theorem 2.1, because $\gamma\leq e^{-r\,T_{i}}\leq 1$, we obtain
$\mathbf{P}\left[\sum_{i=1}^{n}X_{i}\,e^{-r\,T_{i}}>x\,,\;\sum_{j=1}^{n}Y_{j}\,e^{-r\,T_{j}}>y\,,\;N(T)=n\right]=\iint_{\left\{\begin{array}[]{c}0\leq t_{1}\leq\cdots\leq t_{n}\leq T\\ t_{n+1}>T\end{array}\right\}}$ (3.20)
×$\times\mathbf{P}\left[\sum_{i=1}^{n}X_{i}\,e^{-r\,t_{i}}>x\,,\;\sum_{j=1}^{n}Y_{j}\,e^{-r\,t_{j}}>y\right]\,\mathbf{P}[\theta_{1}\in dt_{1},\,\ldots,\,\theta_{n+1}\in dt_{n+1}]$
$\sim\iint_{\left\{\begin{array}[]{c}0\leq t_{1}\leq\cdots\leq t_{n}\leq T\\ t_{n+1}>T\end{array}\right\}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}\,e^{-r\,t_{i}}>x\,,\;Y_{j}\,e^{-r\,t_{j}}>y\right]\right)$ (3.23)
×$\times\mathbf{P}[\theta_{1}\in dt_{1},\,\ldots,\,\theta_{n+1}\in dt_{n+1}]\sim\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}e^{-r\,T_{i}}>x,\;Y_{j}e^{-r\,T_{j}}>y,\;N(T)=n\right]\,,$
@@
$I_{1}(x,\,y;\,T)\sim\sum_{n=1}^{m}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X\,e^{-rT_{i}}>x,\;Y\,e^{-rT_{j}}>y\,,\;N(T)=n\right]=$ (3.24)
$\left(\sum_{n=1}^{\infty}-\sum_{n=m+1}^{\infty}\right)\,\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X\,e^{-rT_{i}}>x,\;Y\,e^{-rT_{j}}>y\,,\;N(T)=n\right]=:\sum_{l=3}^{4}I_{l}(x,\,y;\,T)\,.$
@@
$I_{3}(x,\,y;\,T)=\sum_{i=1}^{\infty}\sum_{n=1}^{\infty}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}\,e^{-rT_{i}}>x,\;Y_{j}\,e^{-rT_{j}}>y\,,\;N(T)=n\right]$ (3.25)
@@
For the first term $I_{31}(x,\,y;\,T)$, taking into account that $\{N(t)\,,\;t\geq 0\}$ represents a renewal function, then $(T_{i}-T_{j})$ is independent of $T_{j}$ and furthermore $(T_{i}-T_{j})\stackrel{{\scriptstyle d}}{{=}}T_{i-j}$, where the equality here, means equal in distribution, hence we find
@@
Now, we estimate $I_{4}(x,\,y;\,T)$. It is easy to see that
@@
$\leq\sum_{n=m+1}^{\infty}\sum_{i=1}^{n}\left(\sum_{i\neq j=1}^{n}+\sum_{i=j=i}\right)\,\mathbf{P}\left[X_{i}\,e^{-r\,T_{1}}>x,\;Y_{j}\,e^{-rT_{1}}>y\,,\;T_{n}\leq T\right]=\sum_{n=m+1}^{\infty}$
$\sum_{i=1}^{n}\left(\sum_{i\neq j=1}^{n}+\sum_{i=j=i}\right)\,\int_{0}^{T}\mathbf{P}\left[X_{i}\,e^{-r\,t}>x,\;Y_{j}\,e^{-r\,t}>y\right]\,\mathbf{P}[N(T-t)\geq n-1]\,\mathbf{P}[T_{1}\in dt]$
$\lesssim(C\vee 1)\,\sum_{n=m+1}^{\infty}[n\,(n-1)+2\,n]\int_{0}^{T}\overline{F}(x\,e^{r\,t})\,\overline{G}(y\,e^{r\,t})\,\mathbf{P}[N(T-t)\geq n-1]\,\mathbf{P}[T_{1}\in dt]$
$\lesssim(C\vee 1)\,\sum_{n=m+1}^{\infty}(n^{2}+n)\,\mathbf{P}[N(T)\geq n-1]\,\int_{0}^{T}\overline{F}(x\,e^{r\,t})\,\overline{G}(y\,e^{r\,t})\,\mathbf{P}[T_{1}\in dt]\,.$
as $x\wedge y\to\infty$. Hence, for any large enough $m$, we can find some small enough $\varepsilon>0$, such that $I_{4}(x,\,y;\,T)\lesssim\varepsilon\,(C\vee 1)\,I(x,\,y,\,F,\,G,\,T)\lesssim\varepsilon\,(C\vee 1)\,I_{3}(x,\,y;\,T)$, as $x\wedge y\to\infty$ and letting $m$ tends to infinity, we get $\varepsilon\downarrow 0$. Thus by relations (3.24), (3.29) and last asymptotic inequality we find
@@
as $x\wedge y\to\infty$. Therefore, by (3), (3) and (3.30), we reach the relation (3.13).   ∎
@@
$=$ $\int_{0}^{\infty}\int_{0}^{\infty}\mathbf{P}\left[D_{r}^{(1)}(T)>x-u\,,\;D_{r}^{(2)}(T)>y-z\right]\,H(du,\,dz)$
$\sim$ $\int_{0}^{\infty}\int_{0}^{\infty}\Delta(x-u,\,y-z\;\,T)\,H(du,\,dz)\,,$
@@
and in last step we used Lemma 3.1. Therefore, due to the fact that $F,\,G\in\mathcal{D}\cap\mathcal{L}\subsetneq\mathcal{L}$, we obtain $\Delta(x-u,\,y-z\;\,T)\sim\Delta(x,\,y\;\,T)$, as $x\wedge y\to\infty$, which in combination of the last relation, by dominated convergence theorem we find
@@ 4. Tail distortion risk measures
In [10] was examined the distortion tail risk measure in a similar background risk model. In [10] are developed asymptotic results for the tail distortion risk measure $TDRM$ of the quantity $\Theta\,{\bf X(w)}$ and the weights $w_{i}$ for $i=1,\,\ldots,\,n$, defined as above, with $\Theta$ to be one-dimensional, and therefore common for all business lines, and independent of ${\bf X}$. Also they assume left-continuous distortion function. In [56] was studied the $TDRM$ for right-continuous distortion function in the $X_{n}(\Theta)$ model, where was permitted multivariate systemic risk factors. In this work was used either $MRV$ structure for the vector ${\bf X}$ or some general enough structures of asymptotic independence for its components, which can provide a direct asymptotic expression. However, the vectors ${\bf\Theta}$ and ${\bf X}$ are still independent.
@@
Obviously, the tail distortion risk measure is a distortion risk measure. Furthermore, if the distortion function is the identical function $g(x)=x$, then the tail distortion risk measure coincides with the conditional tail expectation.
@@
From Assumption 4.1 we obtain ${\bf\Theta\,X}\in MRV(\alpha,\,V,\,\mu_{L})$, therefore ${\bf\Theta\,X(w)}\in\mathcal{R}_{-\alpha}$. From relation (4.2) and the fact that

### v3

@@ 3. Application on bi-dimensional renewal risk model
$N(t)=\sum_{i=1}^{\infty}{\bf 1}_{\{\tau_{i}\leq t\}}\,,$
for any $t\geq 0$, with renewal mean
$\lambda(t)=\mathbf{E}[N(t)]=\sum_{i=i}^{\infty}\mathbf{P}[T_{i}\leq t]\,.$
Indeed, the $\{N(t)\,,\;t\geq 0\}$ represents a homogeneous, renewal process, namely the $\{\theta_{i}\,,\;n\in{\mathbb{N}}\}$, with $\theta_{1}=T_{1}$, $\theta_{i}=T_{i}-T_{i-1}$, for integer $i\geq 2$, representing the inter-arrival times between two successive arrival times, is a sequence of independent and identically distributed, positive random variables.
@@
where * is either ’$\max$’ or ’’, and as $T_{*}$ we consider the
@@
We can observe that in this risk model, except the dependence structures appearing in the conditions on the claims, the two business lines are also dependent through the common renewal process. We note that $C>0$ is the common constant of $SAI$ property, for any $(X,\,Y)$, since these random pairs are identically distributed.
@@
$\leq\sum_{n=m+1}^{\left\lfloor(x\wedge y)/L\right\rfloor}n^{2}\,\mathbf{P}\left[X_{i}>\dfrac{x}{n}\,,\;Y_{j}>\dfrac{y}{n}\right]\,\mathbf{P}[N(T)=n]+\mathbf{P}\left[N(T)>\dfrac{x\wedge y}{L}\right]$
@@
$\lesssim K\,\overline{F}\left(x\right)\,\overline{G}\left(y\right)\,\mathbf{E}\left[[N(T)]^{2\,(p+1)}{\bf 1}_{\{N(T)>m\}}\right]\,,$
@@
$\int_{0}^{T}\mathbf{P}\left[X\,e^{-rt}>x,\;Y\,e^{-rt}>y\right]\,\lambda(dt)\sim C\int_{0}^{T}\overline{F}\left(x\,e^{rt}\right)\,\overline{G}\left(y\,e^{rt}\right)\lambda(dt)\,,$
because $T>0$ if finite. Furthermore,
@@
$\qquad\qquad\leq\dfrac{K}{M}\lim_{m\to\infty}\dfrac{1}{\lambda(T)}\,\mathbf{E}\left[[N(T)]^{2\,(p+1)}{\bf 1}_{\{N(T)>m\}}\right]=0\,,$
where the constant $M>0$, stems from relation (3.16), while in the last step we used [46, Lem. 3.2]. Therefore, it holds
$\lim_{m\to\infty}\limsup_{x\wedge y\to\infty}\dfrac{I_{2}(x,\,y;\,T)}{C\,\int_{0}^{T}\overline{F}\left(x\,e^{rt}\right)\,\overline{G}\left(y\,e^{rt}\right)\lambda(dt)}=0\,.$ (3.17)
Now, we estimate the $I_{1}(x,\,y;\,T)$. At first, through the dominated convergence theorem, by Theorem 2.1, because $0\leq e^{-r\,T_{i}}\leq 1$, we obtain
$\mathbf{P}\left[\sum_{i=1}^{n}X_{i}e^{-r\,T_{i}}>x,\,\sum_{j=1}^{n}Y_{j}e^{-r\,T_{j}}>y,\,N(T)=n\right]=\int\cdots\int_{\left\{\begin{array}[]{c}0\leq t_{1}\leq\cdots\leq t_{n}\leq T\\ t_{n+1}>T\end{array}\right\}}$ (3.20)
×$\times\mathbf{P}\left[\sum_{i=1}^{n}X_{i}\,e^{-r\,t_{i}}>x\,,\;\sum_{j=1}^{n}Y_{j}\,e^{-r\,t_{j}}>y\right]\,\mathbf{P}[T_{1}\in dt_{1},\,\ldots,\,T_{n+1}\in dt_{n+1}]$
$\sim\int\cdots\int_{\left\{\begin{array}[]{c}0\leq t_{1}\leq\cdots\leq t_{n}\leq T\\ t_{n+1}>T\end{array}\right\}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}\,e^{-r\,t_{i}}>x\,,\;Y_{j}\,e^{-r\,t_{j}}>y\right]\right)$ (3.23)
×$\times\mathbf{P}[T_{1}\in dt_{1},\,\ldots,\,T_{n+1}\in dt_{n+1}]\sim\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}e^{-r\,T_{i}}>x,\;Y_{j}e^{-r\,T_{j}}>y,\;N(T)=n\right]\,,$
@@
$I_{1}(x,\,y;\,T)\sim\sum_{n=1}^{m}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}\,e^{-rT_{i}}>x,\;Y_{j}\,e^{-rT_{j}}>y\,,\;N(T)=n\right]=$
$\left(\sum_{n=1}^{\infty}-\sum_{n=m+1}^{\infty}\right)\,\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}\,e^{-rT_{i}}>x,\;Y_{j}\,e^{-rT_{j}}>y\,,\;N(T)=n\right]=:\sum_{l=3}^{4}I_{l}(x,\,y;\,T)\,.$
@@
$I_{3}(x,\,y;\,T)=\sum_{i=1}^{\infty}\sum_{n=i}^{\infty}\sum_{j=1}^{n}\mathbf{P}\left[X_{i}\,e^{-rT_{i}}>x,\;Y_{j}\,e^{-rT_{j}}>y\,,\;N(T)=n\right]$ (3.25)
@@
For the first term $I_{31}(x,\,y;\,T)$, taking into account that $\{N(t)\,,\;t\geq 0\}$ represents a homogeneous renewal process, then $(T_{i}-T_{j})$ is independent of $T_{j}$ and furthermore $(T_{i}-T_{j})\stackrel{{\scriptstyle d}}{{=}}T_{i-j}$, where the equality here, means equal in distribution, hence we find
@@
Now, we estimate $I_{4}(x,\,y;\,T)$. It is easy to see that for sufficiently large $x\wedge y$, we get
@@
$\leq\sum_{n=m+1}^{\infty}\sum_{i=1}^{n}\left(\sum_{i\neq j=1}^{n}+\sum_{i=j=1}^{n}\right)\,\mathbf{P}\left[X_{i}\,e^{-r\,T_{1}}>x,\;Y_{j}\,e^{-rT_{1}}>y\,,\;T_{n}\leq T\right]=\sum_{n=m+1}^{\infty}$
$\sum_{i=1}^{n}\left(\sum_{i\neq j=1}^{n}+\sum_{i=j=1}^{n}\right)\,\int_{0}^{T}\mathbf{P}\left[X_{i}\,e^{-r\,t}>x,\;Y_{j}\,e^{-r\,t}>y\right]\,\mathbf{P}[N(T-t)\geq n-1]\,\mathbf{P}[T_{1}\in dt]$
$\leq K\,\sum_{n=m+1}^{\infty}[n\,(n-1)+2\,n]\int_{0}^{T}\overline{F}(x\,e^{r\,t})\,\overline{G}(y\,e^{r\,t})\,\mathbf{P}[N(T-t)\geq n-1]\,\mathbf{P}[T_{1}\in dt]$
$\leq K\,\sum_{n=m+1}^{\infty}(n^{2}+n)\,\mathbf{P}[N(T)\geq n-1]\,\int_{0}^{T}\overline{F}(x\,e^{r\,t})\,\overline{G}(y\,e^{r\,t})\,\mathbf{P}[T_{1}\in dt]\,.$
where the constant $K>0$, comes from [34, Lem. 3(i)]. Hence, for any large enough $m$ and $x\wedge y>0$, we can find some small enough $\varepsilon>0$, such that $I_{4}(x,\,y;\,T)\leq\varepsilon\,K\,I(x,\,y,\,F,\,G,\,T)\leq\varepsilon\,K\,I_{3}(x,\,y;\,T)$ and letting $m$ tends to infinity, we can take $\varepsilon\downarrow 0$. Thus by relations (3), (3.29) and last asymptotic inequality we find
@@
as $x\wedge y\to\infty$. Therefore, by (3), (3.17) and (3.30), we reach the relation (3.13).   ∎
@@
$=$ $\int_{0}^{\infty}\int_{0}^{\infty}\mathbf{P}\left[D_{r}^{(1)}(T)>x+u\,,\;D_{r}^{(2)}(T)>y+z\right]\,H(du,\,dz)$
$\sim$ $\int_{0}^{\infty}\int_{0}^{\infty}\Delta(x+u,\,y+z\,;\,T)\,H(du,\,dz)\,,$
@@
and in last step we used Lemma 3.1. Therefore, due to the fact that $F,\,G\in\mathcal{D}\cap\mathcal{L}\subsetneq\mathcal{L}$, we obtain $\Delta(x+u,\,y+z\;\,T)\sim\Delta(x,\,y\;\,T)$, as $x\wedge y\to\infty$, which in combination of the last relation, by dominated convergence theorem we find
@@ 4. Tail distortion risk measures
In [10] was examined the distortion tail risk measure in a similar background risk model. In [10] are developed asymptotic results for the tail distortion risk measure $TDRM$ of the quantity $\Theta\,{\bf X(w)}$ and the weights $w_{i}$ for $i=1,\,\ldots,\,n$, defined as above, with $\Theta$ to be one-dimensional, and therefore is a common systemic risk factor for all business lines, and independent of ${\bf X}$. Also they assume left-continuous distortion function. In [56] was studied the $TDRM$ for right-continuous distortion function in the $X_{n}(\Theta)$ model, where was permitted multivariate systemic risk factors. In this work was used either $MRV$ structure for the vector ${\bf X}$ or some general enough structures of asymptotic independence for its components (with regularly varying tails), which can provide a direct asymptotic expression. However, the vectors ${\bf\Theta}$ and ${\bf X}$ are still independent.
@@
It is well-known that, the tail distortion risk measure of a continuous random variable $X$, is a distortion risk measure. Furthermore, if the distortion function is the identical function $g(x)=x$, then the tail distortion risk measure coincides with the conditional tail expectation.
@@
From Assumption 4.1 we obtain ${\bf\Theta\,X}\in MRV(\alpha,\,V,\,\mu)$, therefore ${\bf\Theta\,X(w)}\in\mathcal{R}_{-\alpha}$. From relation (4.2) and the fact that
@@
5. Conclusion
In the corresponding literature, we find a confusion with respect to term multivariate heavy tails. The reason for this, is that the multivariate distribution tail can not be determined uniquely. Having in mind the single big jump principle, the $d$-dimensional, non-weighted form of (2.1) is focused on the following joint distribution tail
$\mathbf{P}\left[\sum_{i=1}^{n_{1}}X_{i}^{(1)}>x_{1},\,\ldots,\,\sum_{i=1}^{n_{d}}X_{i}^{(d)}>x_{d}\right]\sim\sum_{i=1}^{n_{1}}\cdots\sum_{i=1}^{n_{d}}\mathbf{P}\left[X_{i}^{(1)}>x_{1},\,\ldots,\,X_{i}^{(d)}>x_{d}\right]\,,$ (5.1)
as $\wedge_{i=1}^{d}x_{i}\to\infty$. Along with approach of the multivariate subexponentiality by [41], we find in [29], under different conditions the relation
$\mathbf{P}\left[\sum_{i=1}^{n}{\bf X}^{(i)}\in x\,A\right]\sim\sum_{i=1}^{n}\mathbf{P}\left[{\bf X}^{(i)}\in x\,A\right]\,,$ (5.2)
as $x\to\infty$, where ${\bf X}^{(1)},\,\ldots,\,{\bf X}^{(n)}$ represent $d$-dimensional random vectors, and $A$ is some rare set. Although the single big jump approximation in (5.2) has several good properties, as for example that most of the closure properties of the uni-variate distribution classes hold also for the multivariate ones and further some results can be generalized to multidimensional set up, still there exist some basic drawbacks.
The main drawback is that the set $A$ does not represent the joint distribution tail, see in [29, Rem. 2.2]. This leads to a ’linear’ approximation of the single big jump, in the sense that in order to happen this representation, should be the from $1$ up to $d$, from the total $n\times d$, random variables, sufficiently ’large’. As consequence, the approximation of relation (5.2) can not represent the probabilities $\psi_{\max}$ and $\psi$, belonging among the four most popular ruin probabilities in multivariate risk models.
In opposite direction, relation (5.1), studied in section 2 with $d=2$, can solve the problem with these two ruin probabilities, since it is focused on the joint distribution tail. Furthermore, we can say that relation (5.1), follows a ’non-linear’ approach of the single big jump, since requires $d$ sufficiently ’large’ random variables, namely one big jump for each line of business, if you look from the risk theory aspect. In this sense, the two approached by (5.1) and (5.2), work in complementary modes.
Finally, we want to mention that the $MRV$, although can give solutions with respect to ’$\psi_{\max}$’ and ’$\psi_{and}$’, satisfies only the linear single big jump, namely relation (5.2). From this point of view the study of (5.1) leads to revision of multivariate heavy-tailed distributions, even of the well-established ones, like the $MRV$. The events described by relation (5.2) do NOT put emphasis in the dimension, in opposite with the events described by relation (5.1).
Acknowledgments. We feel the need to express our gratitude to two anonymous referees, for their crucial remarks, that improved the present paper.